OpenStudy (anonymous):

Answer this for me Please and Thank you, Square Roots and Irrational Numbers!

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt 64 = \sqrt 16 * \sqrt 16 = \sqrt 4 * \sqrt 4 * \sqrt 4 * \sqrt 4 = 2 * 2 * 2 * 2 = 16\]

OpenStudy (anonymous):

Woah this is way off

OpenStudy (anonymous):

\[\sqrt 64 = \sqrt 16 * \sqrt 4 = 2 * \sqrt 4 * \sqrt 4 = 2 * 2 * 2\]

OpenStudy (anonymous):

= 8